OpenStudy (anonymous):

A company did a quality check on all the packs of trail mix it manufactured. Each pack of trail mix is targeted to weigh 9.25 oz. A pack must weigh within 0.23 oz of the target weight to be accepted. What is the range of rejected masses, x, for the manufactured trail mix? x < 9.02 or x > 9.48 because |x - 0.23| + 9.25 > 0 x < 9.25 or x > 9.48 because |x - 9.25| > 0.23 x < 9.25 or x > 9.48 because |x - 0.23| + 9.25 > 0 x < 9.02 or x > 9.48 because |x - 9.25| > 0.23

OpenStudy (anonymous):

@jigglypuff314 @phi @ganeshie8

jigglypuff314 (jigglypuff314):

so it must be within 9.25 - 0.23 = ? ans 9.25 + 0.23 = ?

OpenStudy (anonymous):

9.25 - 0.23 = 9.02 9.25 + 0.23 = 9.48

jigglypuff314 (jigglypuff314):

yes :) then try solving the absolute values do you know how to do that?

OpenStudy (anonymous):

|x - 9.25| |x - 0.23| I don't really know how to, the text book did a terrible job of explaining... I kept reviewing and I still don't understand.

jigglypuff314 (jigglypuff314):

it's alright :) when given |a - b| > c it can be separated into a - b > c and a - b < -c do you understand it like that? or would you like for my to give an example? :)

OpenStudy (anonymous):

Um, i think i get it, im thinking the answer is B but i am not sure,

jigglypuff314 (jigglypuff314):

close, you're correct on the absolute values parts but what were the numbers I had you get?

OpenStudy (anonymous):

9.02 and 9.48

OpenStudy (anonymous):

Ahh wait I think I get it! It's D right??

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

Thanks! Can you help with a couple more?

jigglypuff314 (jigglypuff314):

sure :) post them as separate question posts though :P

OpenStudy (anonymous):

Okay, thank you so much

OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality 3x + 1 > 10 and 4x less than or greater to 20. A number line with an open circle on 3, a closed circle on 5, and shading in between A number line with an open circle on 3, shading to the left, and a closed circle on 5, shading to the right A number line with a closed circle on 3, shading to the left, and an open circle on 5, and shading to the right A number line with a closed circle on 3, an open circle on 5, and shading in between

jigglypuff314 (jigglypuff314):

solve for x from 3x + 1 > 10 can you do that? it's just like any other equation :)

OpenStudy (anonymous):

x>3

jigglypuff314 (jigglypuff314):

correct :) then 4x ≤ 20 solve for x

OpenStudy (anonymous):

x ≤ 5

jigglypuff314 (jigglypuff314):

correct :) good job! now, do you know how to graph those two on the number line?

OpenStudy (anonymous):

I'm going to go with A and yes I do

jigglypuff314 (jigglypuff314):

A is correct :)

OpenStudy (anonymous):

thanks! :) I have just two more, I might bother you again later on today though, i like math yet I suck at it, lol

OpenStudy (anonymous):

Jess observed that 60% of the people at a mall on a particular day shopped for clothes. If 2,500 people at the mall did not shop for clothes that day, the number of people who shopped for clothes that day was ______. (only put numeric values, no other symbols)

jigglypuff314 (jigglypuff314):

100% - 60% = 40% 40% = 2500 if x = total 0.4 x = 2500 solve for x then multiply x by 0.6

OpenStudy (anonymous):

x = 6250 6250 * 0.6 = 3750

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (anonymous):

okay, is that it for this question?

jigglypuff314 (jigglypuff314):

I think so :P

OpenStudy (anonymous):

okay, i have one more, one minute

OpenStudy (anonymous):

The equation below shows the area of a trapezoid, A, with a height of 9 cm, and one base 35 cm. : A = 9 over 2(b + 35) Which of the following formulas correctly solves for the other base, b? b = 2A over 9 + 35 b = 2 multiplied by A over 9 - 35 b = 2 multiplied by A plus 35, all over 9 b = 2 multiplied by A minus 35, all over 9

jigglypuff314 (jigglypuff314):

to solve for b multiply both sides by 2 divide both sides by 9 subtract 35 from both sides

OpenStudy (anonymous):

so it's D?

jigglypuff314 (jigglypuff314):

I did divide by 9 before subtracting 35 so it won't be "all over 9"

OpenStudy (anonymous):

Oh ok, so it will be B then

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (anonymous):

okay, thank you so much, i wish i could give you five more medals, :)

jigglypuff314 (jigglypuff314):

lol no need ;) I have more than enough medals :P